OpenStudy (anonymous):

I GIVE MEDALS Solve: 6(x -2) + 3(x -2) = 18 A. x = 2 B. x = 22/9 C. x =4

OpenStudy (solomonzelman):

6(x -2) + 3(x -2) = 18 9(x -2) = 18 x-2=2 x=?

OpenStudy (anonymous):

6(x-2) + 3(x-2) = 18 Multiply out brackets. So, 6x - 12 + 3x - 6 = 18 Now you can just reorder the terms and simplify to solve for your variable. REORDER: 6x+ 3x - 12 - 6 = 18 9x - 18 = 18 9x - 18 + 18 = 18 +18 9x = 36 Divide by 9 on both sides. x = 4

OpenStudy (anonymous):

@sarara thank you so much omfg